# PrivTune: Efficient and Privacy-Preserving Fine-Tuning of Large Language Models via Device-Cloud Collaboration

### v2

@@ I Introduction
The widespread deployment of Large Language Models (LLMs) in mobile and edge applications has spawned Language Model-as-a-Service (LMaaS), with pre-trained LLMs cloud-hosted for users to submit task queries [sun2022black, thapa2022splitfed, openai_finetuning_guide, su2024titanic]. As task complexity grows, users increasingly require customized LLMs (e.g., personalized intelligent assistants) to cope with personalized scenarios [cai2024llmaas]. To meet this demand, service providers have introduced fine-tuning APIs that allow users to customize models on their private data. For example, OpenAI’s fine-tuning API enables users to upload domain-specific data (e.g., personal documents) in JSON format for tailored model development [openai_finetuning_guide]. However, this approach requires uploading sensitive data to the cloud, creating a fundamental tension between performance and privacy. Although platforms like OpenAI claim to delete user data within 30 days [openai_finetuning_guide], such assurances may not be sufficient to alleviate concerns among privacy-sensitive users. This privacy concern is exacerbated in scenarios where data is inherently sensitive and regulated, e.g., in healthcare [arora2023promise] and finance [xu2024fwdllm].
@@
To alleviate such privacy concerns, a straightforward approach is the on-device adaptation technique [zhuang2024litemoe, ding2023parameter, gao2025federated]. The key idea is to allow user clients to download LLMs from the cloud and perform local fine-tuning by updating only a small set of trainable parameters, i.e., using Parameter-Efficient Fine-Tuning (PEFT) [zhuang2024litemoe, yu2024edge, kuang2024federatedscope] methods such as Low-Rank Adaptation (LoRA) [hulora], without uploading sensitive data. However, these approaches typically assume full access to the cloud LLM weights on the device, and thus such deployment may violate model ownership restrictions  [zhang2025secret, wang2025never, duan2023flocks, wangflora].
@@
In light of the above challenges, device-cloud collaboration techniques are developed to achieve secure and efficient fine-tuning. These techniques propose offloading part of the fine-tuning task to the device, enabling the cloud and device to collaboratively complete fine-tuning of LLMs [xiao2023offsite, yao2025scaleot, wu2024fedbiot, yao2024split, lin2024splitlora, li2023privacy]. Specifically, the cloud offloads a portion of the LLM (e.g., the embedding layers) [yao2024split, lin2024splitlora, shen2025sap] or a compressed version of the model [xiao2023offsite, yao2025scaleot, wu2024fedbiot], to the user device. The user then uses this component to generate intermediate representations (e.g., embeddings or model updates), which are shared with the cloud for joint fine-tuning. To further protect the privacy of these representations, one naive way is to instruct devices to inject Differential Privacy (DP) [dwork2006differential] noise before transmission, mitigating the risk of privacy inference attacks (e.g., embedding inversion attacks) [zhu2019deep, qu2025prompt, song2020information, huang2024transferable]. In this way, device-cloud collaboration avoids exposing the full cloud LLM to edge devices, reduces local resource demands, and keeps raw data on the device. However, this paradigm has a critical bottleneck in practice: the privacy-utility trade-off is difficult to achieve due to the injection of costly DP noise.
@@
To this end, we formulate the noise design as an optimization problem that jointly considers both defense and utility objectives. Specifically, we aim to minimize the similarity between semantically related $n$-hop indirect neighbor tokens and maximize the intra-class cluster distance. This disrupts structured token distributions and enhances resistance to privacy inference attacks. Building atop, we further constrain the $\ell_{2}$-norm and distribution drift of embeddings before and after perturbation to preserve model utility. However, we notice that while the optimization formulation yields an optimal noise vector, it is deterministic. This makes it vulnerable to reverse engineering by an attacker, as the privacy guarantee fundamentally relies on noise being a random perturbation rather than a predictable signal [dwork2006differential, pan2020privacy, feyisetan2020privacy]. To resolve this, we design a token importance-aware $d_{\chi}$}-Privacy mechanism. It uses the optimal noise vector to set the mean of the noise distribution, ensuring that the expected noise direction aligns with the optimization goal and meets randomness. Finally, we compute token importance scores using statistical features and attention weights, and use them to guide weighted noise injection, i.e., adding more noise to less important tokens while protecting critical semantic information.
We conduct extensive experiments on five benchmark datasets, covering both classification and generation tasks, and evaluate PrivTune across LLMs of varying sizes under three embedding inversion attacks and three attribute inference attacks. Specifically, we compare PrivTune with five state-of-the-art (SOTA) baselines, i.e., DP-Forward [du2023dp], SLFT-DP [WANG2024Selective], SAP [shen2025sap], ScaleOT [yao2025scaleot], and FedBiOT [wu2024fedbiot]. Our results demonstrate that PrivTune effectively mitigates these privacy attacks while preserving model utility. Specifically, PrivTune reduces the success rate of embedding inversion attacks to approximately 10%, with only a 3.33% drop in utility performance, surpassing all existing SOTA baselines. Moreover, experimental results show that PrivTune is efficient in terms of computational and communication overhead, highlighting its effectiveness in real-world scenarios.
@@ II Related Work
PEFT methods have emerged as a practical solution to reduce the computational and memory overhead of adapting LLMs [hulora, ding2023parameter, lester2021power] on device. Instead of updating the entire model, PEFT focuses on fine-tuning a small subset of parameters, enabling faster and more resource-efficient adaptation. For example, LoRA [hulora] introduces low-rank matrices into existing weight matrices, significantly reducing the number of parameters that need to be updated. Despite their efficiency, most PEFT approaches assume full access to the LLM and lack built-in privacy protections, making them unsuitable for sensitive on-device applications where user data privacy must be preserved [xu2024fwdllm].
@@
The growing demand for data privacy and personalization has driven research on device–cloud collaboration for fine-tuning of LLMs. Existing research has made preliminary progress along two main directions: Offsite-Tuning [xiao2023offsite, yao2025scaleot, wu2024fedbiot] and Split Learning-based Fine-Tuning (SLFT) [yao2024split, lin2024splitlora, li2023privacy]. To start with, Offsite-Tuning avoids sharing the full cloud LLM to edge devices. Instead, it transmits a compressed version of the model, referred to as an Emulator, using model compression [xiao2023offsite] or knowledge distillation [yao2025scaleot] techniques, enabling local fine-tuning via on-device adapters (e.g., LoRA). While this approach mitigates privacy concerns for users (as now the fine-tuning process will happen locally on the users’ devices), it still poses risks to model ownership due to the structural similarity between the Emulator and the original cloud LLM [wu2024fedbiot, shen2025sap]. In contrast, SLFT offloads part of the cloud LLM (called the bottom model) to the edge device and enables privacy-preserving fine-tuning by exchanging differentially private intermediate embeddings [li2023privacy]. While this makes it harder for the edge device to reconstruct the full model, the added noise from DP [dwork2006differential] often leads to a poor privacy-utility trade-off.
@@ III Background
We consider an SLFT scenario where a service provider $\mathcal{S}$ holds an LLM parameterized by $\bm{w}$, and offers customized fine-tuning services to users $\mathcal{U}$, where each user $u_{i}$ possesses a private dataset $\mathcal{D}_{i}$ containing sensitive textual data. To enable secure and resource-efficient fine-tuning, the LLM is partitioned into two components via SL [thapa2022splitfed, romanini2021pyvertical]: a bottom model $f_{\theta_{c}}$ and a top model $g_{\theta_{s}}$, such that the full model is represented as: $\mathcal{F}(x;\bm{w})=g_{\theta_{s}}(f_{\theta_{c}}(x))$, where $f_{\theta_{c}}$ is deployed locally on the user’s edge device, $g_{\theta_{s}}$ is retained on the server, $\bm{w}=\theta_{c}\cup\theta_{s}$ denotes the overall model parameters. The forward pass proceeds as follows. Given an input $x\in\mathcal{D}_{i}$, the device computes an intermediate representation: $\bm{h}=f_{\theta_{c}}(x),$ and transmits $\bm{h}$ to the server. The server then computes the final prediction: $\hat{y}=g_{\theta_{s}}(\bm{h}).$ Since the label $y$ remains private and is only available on the user side, the backward pass is collaboratively performed. The above process can be formulated as:
@@
where $\bm{\delta}$ is the trainable parameters of the PEFT methods, e.g., LoRA [hulora]. This training process proceeds iteratively until fine-tuned model convergence. To prevent SLFT from privacy inference attacks [du2023dp, song2020information, shokri2017membership, huang2024transferable] (detailed in the following sections), existing approaches typically rely on DP noise mechanisms to provide formal privacy guarantees. However, such methods often incur significant performance degradation. This highlights the need for slight perturbation-based defenses that can offer privacy protection with minimal impact on utility, as explored in later sections (see §V-D).
@@
Attack-0: Direct Activation Inversion [song2020information]. Given the activation $\bm{h}=\phi_{d}(x)$, an attacker attempts to find $\hat{x}$ such that:
@@
Attack-1: Gradient Inversion Attack [zhu2019deep]. When gradients $g=\nabla_{\theta_{c}}\mathcal{L}(f_{\theta_{c}}(x),y)$ are visible to an attacker (e.g., via model updates), the attack recovers input $x$ by minimizing:
@@
Attack-2: Embedding Matching and Nearest Neighbor Recovery [huang2024transferable]. In some cases, attackers can simply compute cosine similarity between leaked embeddings $\bm{h}_{i}$ and all entries in the vocabulary embedding matrix $\mathcal{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$:
@@
Attack-3: Supervised Attribute Classifier [du2023dp]. With access to a labeled shadow dataset, the attacker trains a dedicated classifier $f_{\theta}$ to predict private attributes directly from intermediate features. Specifically, the objective of the attacker is defined as follows:
@@
Attack-4: Gradient-based Attribute Inference [shokri2017membership]. In multi-round SL, when the server can observe gradients $\nabla_{\theta_{s}}\mathcal{L}$, the attacker hypothesizes that certain attributes disproportionately influence these updates. The attacker computes:
@@
Attack-5: Representation Clustering Attack [song2020information]. The attacker applies unsupervised clustering (e.g., K-means) to the intermediate activations $\bm{h}$ collected during SL. After unsupervised clustering, the attacker maps cluster centroids to attribute classes using shadow-labeled data. The core assumption is:
@@ IV-A Threat Model
Since our work focuses on developing defenses, we assume a strong adversarial model where the attacker possesses significant background knowledge. Specifically, the attacker is assumed to have access to the intermediate representations transmitted by the user, as well as full knowledge of the user’s bottom model, including its parameters, architecture, and version, i.e., a white-box setting. Additionally, the attacker is assumed to have white-box access to the perturbed intermediate representations produced by our defense mechanism. However, the attacker does not know the specific perturbation technique applied, as it is carefully designed and kept confidential by the defender [lou2025grid].
@@
We assume the attacker is capable of arbitrarily manipulating these intermediate representations. It then uses them to infer private information by adopting the state-of-the-art embedding inversion attacks [du2023dp] and attribute inference attacks [song2020information]. However, the attacker cannot alter the design or execution of the fine-tuning service protocol or the defense mechanisms employed by the edge device.
@@ IV-B Defense Formulation
Given the above threat model, aligned with the previous work [wang2023privatelora, qu2021natural], our goal is to achieve: 1) the service provider cannot recover the original input text or infer private attributes from the transmitted text representation; 2) the proposed defense should maintain comparable performance compared to non-private methods. Next, we detail our motivation and provide mathematical formulations for our defense methods.
@@
In doing so, we make it significantly harder for the attacker to invert or infer private information from intermediate outputs [huang2024transferable, abadi2016deep]. Additionally, we will formulate this noise generation as an optimization problem to ensure that both defense goals and model utility are maintained.
@@
Semantic Preservation. We constrain the angular deviation between the original and perturbed embeddings to ensure semantic information is preserved [qiu2025flm]:
@@ V-A Overview
In this paper, we propose PrivTune, a device-cloud collaboration framework built upon the SplitNN architecture (the classic SL configuration) [thapa2022splitfed, romanini2021pyvertical] to enable privacy-preserving and efficient fine-tuning of LLMs, as illustrated in Fig. 1. The framework consists of three key stages: model partition, on-device perturbation, and collaborative training. Next, we briefly introduce the execution details of each stage.
@@
At this stage, the edge device performs a forward pass through its local embedding layers to generate intermediate token representations. To reduce computational overhead on resource-constrained devices, we adopt the post-fine-tuning approach [sun2022black], wherein the embedding layers and Transformer modules of the bottom model are frozen during fine-tuning. We first compute an optimal noise vector by solving OPT-2 (see §V-C). This vector serves as a directional guide for token importance-aware $d_{\chi}$}-Privacy noise sampling (see §V-D), where the sampled noise is then added to the intermediate representation to protect privacy while preserving model utility.
@@ V-C2 Reformulating Constraints
This constraint implicitly preserves both model prediction behavior (assuming local Lipschitz continuity of $g_{\theta_{s}}$ [hulora]) and the semantic integrity of the input.
@@ V-C3 Reformulation of OPT-2
To efficiently solve OPT-3, we employ Projected Gradient Descent (PGD) [madry2018towards] with dual projection steps to enforce the two convex constraints. The optimization is performed over the perturbation vector $\bm{p}_{i}$, where the perturbed embedding is defined as $\tilde{\bm{h}}_{i}=\bm{h}_{i}+\bm{p}_{i}$. At each iteration, we compute the gradient of the objective $\mathcal{L}=D_{i}^{\mathrm{EIA}}-D_{i}^{\mathrm{AIA}}$ with respect to $\bm{p}_{i}$, perform a gradient descent update, and project the result back onto the feasible region defined by the local perturbation constraint (L1) and the global distributional constraint (L2). Specifically, if the perturbation exceeds the local bound $\|\tilde{\bm{h}}_{i}-\bm{h}_{i}\|_{2}^{2}>2B^{2}(1-\delta)$, we scale it back to lie on the boundary of the $\ell_{2}$-ball centered at $\bm{h}_{i}$; similarly, if $\|\tilde{\bm{h}}_{i}-\bm{\mu}\|_{2}>R$, we project $\tilde{\bm{h}}_{i}$ back toward the global centroid $\bm{\mu}$ with radius $R$. We denote the optimal noise vector solved as $\bm{p}_{i}^{*}$. The procedure is efficient, parallelizable, and well-suited for on-device deployment, with convergence typically achieved in a small number of iterations when the learning rate is properly tuned. The details can be found in Algo. 1.

### v3

@@ I Introduction
The widespread deployment of Large Language Models (LLMs) in mobile and edge applications has spawned Language Model-as-a-Service (LMaaS), with pre-trained LLMs cloud-hosted for users to submit task queries [36, 37, 26, 35]. As task complexity grows, users increasingly require customized LLMs (e.g., personalized intelligent assistants) to cope with personalized scenarios [5]. To meet this demand, service providers have introduced fine-tuning APIs that allow users to customize models on their private data. For example, OpenAI’s fine-tuning API enables users to upload domain-specific data (e.g., personal documents) in JSON format for tailored model development [26]. However, this approach requires uploading sensitive data to the cloud, creating a fundamental tension between performance and privacy. Although platforms like OpenAI claim to delete user data within 30 days [26], such assurances may not be sufficient to alleviate concerns among privacy-sensitive users. This privacy concern is exacerbated in scenarios where data is inherently sensitive and regulated, e.g., in healthcare [4] and finance [45].
@@
To alleviate such privacy concerns, a straightforward approach is the on-device adaptation technique [52, 6, 11]. The key idea is to allow user clients to download LLMs from the cloud and perform local fine-tuning by updating only a small set of trainable parameters, i.e., using Parameter-Efficient Fine-Tuning (PEFT) [52, 49, 16] methods such as Low-Rank Adaptation (LoRA) [13], without uploading sensitive data. However, these approaches typically assume full access to the cloud LLM weights on the device, and thus such deployment may violate model ownership restrictions  [50, 39, 8, 42].
@@
In light of the above challenges, device-cloud collaboration techniques are developed to achieve secure and efficient fine-tuning. These techniques propose offloading part of the fine-tuning task to the device, enabling the cloud and device to collaboratively complete fine-tuning of LLMs [44, 48, 43, 47, 20, 18]. Specifically, the cloud offloads a portion of the LLM (e.g., the embedding layers) [47, 20, 32] or a compressed version of the model [44, 48, 43], to the user device. The user then uses this component to generate intermediate representations (e.g., embeddings or model updates), which are shared with the cloud for joint fine-tuning. To further protect the privacy of these representations, one naive way is to instruct devices to inject Differential Privacy (DP) [9] noise before transmission, mitigating the risk of privacy inference attacks (e.g., embedding inversion attacks) [51, 30, 34, 14]. In this way, device-cloud collaboration avoids exposing the full cloud LLM to edge devices, reduces local resource demands, and keeps raw data on the device. However, this paradigm has a critical bottleneck in practice: the privacy-utility trade-off is difficult to achieve due to the injection of costly DP noise.
@@
To this end, we formulate the noise design as an optimization problem that jointly considers both defense and utility objectives. Specifically, we aim to minimize the similarity between semantically related $n$-hop indirect neighbor tokens and maximize the intra-class cluster distance. This disrupts structured token distributions and enhances resistance to privacy inference attacks. Building atop, we further constrain the $\ell_{2}$-norm and distribution drift of embeddings before and after perturbation to preserve model utility. However, we notice that while the optimization formulation yields an optimal noise vector, it is deterministic. This makes it vulnerable to reverse engineering by an attacker, as the privacy guarantee fundamentally relies on noise being a random perturbation rather than a predictable signal [9, 27, 10]. To resolve this, we design a token importance-aware $d_{\chi}$}-Privacy mechanism. It uses the optimal noise vector to set the mean of the noise distribution, ensuring that the expected noise direction aligns with the optimization goal and meets randomness. Finally, we compute token importance scores using statistical features and attention weights, and use them to guide weighted noise injection, i.e., adding more noise to less important tokens while protecting critical semantic information.
We conduct extensive experiments on five benchmark datasets, covering both classification and generation tasks, and evaluate PrivTune across LLMs of varying sizes under three embedding inversion attacks and three attribute inference attacks. Specifically, we compare PrivTune with five state-of-the-art (SOTA) baselines, i.e., DP-Forward [7], SLFT-DP [40], SAP [32], ScaleOT [48], and FedBiOT [43]. Our results demonstrate that PrivTune effectively mitigates these privacy attacks while preserving model utility. Specifically, PrivTune reduces the success rate of embedding inversion attacks to approximately 10%, with only a 3.33% drop in utility performance, surpassing all existing SOTA baselines. Moreover, experimental results show that PrivTune is efficient in terms of computational and communication overhead, highlighting its effectiveness in real-world scenarios.
@@ II Related Work
PEFT methods have emerged as a practical solution to reduce the computational and memory overhead of adapting LLMs [13, 6, 17] on device. Instead of updating the entire model, PEFT focuses on fine-tuning a small subset of parameters, enabling faster and more resource-efficient adaptation. For example, LoRA [13] introduces low-rank matrices into existing weight matrices, significantly reducing the number of parameters that need to be updated. Despite their efficiency, most PEFT approaches assume full access to the LLM and lack built-in privacy protections, making them unsuitable for sensitive on-device applications where user data privacy must be preserved [45].
@@
The growing demand for data privacy and personalization has driven research on device–cloud collaboration for fine-tuning of LLMs. Existing research has made preliminary progress along two main directions: Offsite-Tuning [44, 48, 43] and Split Learning-based Fine-Tuning (SLFT) [47, 20, 18]. To start with, Offsite-Tuning avoids sharing the full cloud LLM to edge devices. Instead, it transmits a compressed version of the model, referred to as an Emulator, using model compression [44] or knowledge distillation [48] techniques, enabling local fine-tuning via on-device adapters (e.g., LoRA). While this approach mitigates privacy concerns for users (as now the fine-tuning process will happen locally on the users’ devices), it still poses risks to model ownership due to the structural similarity between the Emulator and the original cloud LLM [43, 32]. In contrast, SLFT offloads part of the cloud LLM (called the bottom model) to the edge device and enables privacy-preserving fine-tuning by exchanging differentially private intermediate embeddings [18]. While this makes it harder for the edge device to reconstruct the full model, the added noise from DP [9] often leads to a poor privacy-utility trade-off.
@@ III Background
We consider an SLFT scenario where a service provider $\mathcal{S}$ holds an LLM parameterized by $\bm{w}$, and offers customized fine-tuning services to users $\mathcal{U}$, where each user $u_{i}$ possesses a private dataset $\mathcal{D}_{i}$ containing sensitive textual data. To enable secure and resource-efficient fine-tuning, the LLM is partitioned into two components via SL [37, 31]: a bottom model $f_{\theta_{c}}$ and a top model $g_{\theta_{s}}$, such that the full model is represented as: $\mathcal{F}(x;\bm{w})=g_{\theta_{s}}(f_{\theta_{c}}(x))$, where $f_{\theta_{c}}$ is deployed locally on the user’s edge device, $g_{\theta_{s}}$ is retained on the server, $\bm{w}=\theta_{c}\cup\theta_{s}$ denotes the overall model parameters. The forward pass proceeds as follows. Given an input $x\in\mathcal{D}_{i}$, the device computes an intermediate representation: $\bm{h}=f_{\theta_{c}}(x),$ and transmits $\bm{h}$ to the server. The server then computes the final prediction: $\hat{y}=g_{\theta_{s}}(\bm{h}).$ Since the label $y$ remains private and is only available on the user side, the backward pass is collaboratively performed. The above process can be formulated as:
@@
where $\bm{\delta}$ is the trainable parameters of the PEFT methods, e.g., LoRA [13]. This training process proceeds iteratively until fine-tuned model convergence. To prevent SLFT from privacy inference attacks [7, 34, 33, 14] (detailed in the following sections), existing approaches typically rely on DP noise mechanisms to provide formal privacy guarantees. However, such methods often incur significant performance degradation. This highlights the need for slight perturbation-based defenses that can offer privacy protection with minimal impact on utility, as explored in later sections (see §V-D).
@@
Attack-0: Direct Activation Inversion [34]. Given the activation $\bm{h}=\phi_{d}(x)$, an attacker attempts to find $\hat{x}$ such that:
@@
Attack-1: Gradient Inversion Attack [51]. When gradients $g=\nabla_{\theta_{c}}\mathcal{L}(f_{\theta_{c}}(x),y)$ are visible to an attacker (e.g., via model updates), the attack recovers input $x$ by minimizing:
@@
Attack-2: Embedding Matching and Nearest Neighbor Recovery [14]. In some cases, attackers can simply compute cosine similarity between leaked embeddings $\bm{h}_{i}$ and all entries in the vocabulary embedding matrix $\mathcal{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$:
@@
Attack-3: Supervised Attribute Classifier [7]. With access to a labeled shadow dataset, the attacker trains a dedicated classifier $f_{\theta}$ to predict private attributes directly from intermediate features. Specifically, the objective of the attacker is defined as follows:
@@
Attack-4: Gradient-based Attribute Inference [33]. In multi-round SL, when the server can observe gradients $\nabla_{\theta_{s}}\mathcal{L}$, the attacker hypothesizes that certain attributes disproportionately influence these updates. The attacker computes:
@@
Attack-5: Representation Clustering Attack [34]. The attacker applies unsupervised clustering (e.g., K-means) to the intermediate activations $\bm{h}$ collected during SL. After unsupervised clustering, the attacker maps cluster centroids to attribute classes using shadow-labeled data. The core assumption is:
@@ IV-A Threat Model
Since our work focuses on developing defenses, we assume a strong adversarial model where the attacker possesses significant background knowledge. Specifically, the attacker is assumed to have access to the intermediate representations transmitted by the user, as well as full knowledge of the user’s bottom model, including its parameters, architecture, and version, i.e., a white-box setting. Additionally, the attacker is assumed to have white-box access to the perturbed intermediate representations produced by our defense mechanism. However, the attacker does not know the specific perturbation technique applied, as it is carefully designed and kept confidential by the defender [22].
@@
We assume the attacker is capable of arbitrarily manipulating these intermediate representations. It then uses them to infer private information by adopting the state-of-the-art embedding inversion attacks [7] and attribute inference attacks [34]. However, the attacker cannot alter the design or execution of the fine-tuning service protocol or the defense mechanisms employed by the edge device.
@@ IV-B Defense Formulation
Given the above threat model, aligned with the previous work [41, 29], our goal is to achieve: 1) the service provider cannot recover the original input text or infer private attributes from the transmitted text representation; 2) the proposed defense should maintain comparable performance compared to non-private methods. Next, we detail our motivation and provide mathematical formulations for our defense methods.
@@
In doing so, we make it significantly harder for the attacker to invert or infer private information from intermediate outputs [14, 1]. Additionally, we will formulate this noise generation as an optimization problem to ensure that both defense goals and model utility are maintained.
@@
Semantic Preservation. We constrain the angular deviation between the original and perturbed embeddings to ensure semantic information is preserved [28]:
@@ V-A Overview
In this paper, we propose PrivTune, a device-cloud collaboration framework built upon the SplitNN architecture (the classic SL configuration) [37, 31] to enable privacy-preserving and efficient fine-tuning of LLMs, as illustrated in Fig. 1. The framework consists of three key stages: model partition, on-device perturbation, and collaborative training. Next, we briefly introduce the execution details of each stage.
@@
At this stage, the edge device performs a forward pass through its local embedding layers to generate intermediate token representations. To reduce computational overhead on resource-constrained devices, we adopt the post-fine-tuning approach [36], wherein the embedding layers and Transformer modules of the bottom model are frozen during fine-tuning. We first compute an optimal noise vector by solving OPT-2 (see §V-C). This vector serves as a directional guide for token importance-aware $d_{\chi}$}-Privacy noise sampling (see §V-D), where the sampled noise is then added to the intermediate representation to protect privacy while preserving model utility.
@@ V-C2 Reformulating Constraints
This constraint implicitly preserves both model prediction behavior (assuming local Lipschitz continuity of $g_{\theta_{s}}$ [13]) and the semantic integrity of the input.
@@ V-C3 Reformulation of OPT-2
To efficiently solve OPT-3, we employ Projected Gradient Descent (PGD) [24] with dual projection steps to enforce the two convex constraints. The optimization is performed over the perturbation vector $\bm{p}_{i}$, where the perturbed embedding is defined as $\tilde{\bm{h}}_{i}=\bm{h}_{i}+\bm{p}_{i}$. At each iteration, we compute the gradient of the objective $\mathcal{L}=D_{i}^{\mathrm{EIA}}-D_{i}^{\mathrm{AIA}}$ with respect to $\bm{p}_{i}$, perform a gradient descent update, and project the result back onto the feasible region defined by the local perturbation constraint (L1) and the global distributional constraint (L2). Specifically, if the perturbation exceeds the local bound $\|\tilde{\bm{h}}_{i}-\bm{h}_{i}\|_{2}^{2}>2B^{2}(1-\delta)$, we scale it back to lie on the boundary of the $\ell_{2}$-ball centered at $\bm{h}_{i}$; similarly, if $\|\tilde{\bm{h}}_{i}-\bm{\mu}\|_{2}>R$, we project $\tilde{\bm{h}}_{i}$ back toward the global centroid $\bm{\mu}$ with radius $R$. We denote the optimal noise vector solved as $\bm{p}_{i}^{*}$. The procedure is efficient, parallelizable, and well-suited for on-device deployment, with convergence typically achieved in a small number of iterations when the learning rate is properly tuned. The details can be found in Algo. 1.